# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v3

@@ IV Conclusion
In this study, we developed and analyzed two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW for improved cell clustering performance on scRNA-seq data. Firstly, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) we developed extends the edge-dependent vertex weight (EDVW) hypergraph random walk proposed by [41], and consistently outperformed it in our experiments. DIPHW’s key idea lies in its symmetrized account of importance: it considers both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering. Secondly, the CoMem and CoMem-DIPHW clustering methods we developed incorporate a memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks to integrate both local and global information when computing embeddings for clustering. Thirdly, we identified and showed with experiments 3 the issue of inflated signals in co-expression networks brought by sparsity of scRNA-seq data, which provides additional motivation for the use of hypergraph representations. Fourthly, we compared our proposed methods with well-established clustering methods commonly used in scRNA-seq analysis packages. Our clustering performance evaluation spanned a broad range of modularity conditions, where CoMem-DIPHW and DIPHW consistently exhibited superior performance, particularly in weak modularity regimes. When clusters were harder to detect, baseline methods experienced a much more pronounced decay in performance compared to CoMem-DIPHW. For empirical analysis, we tested our algorithms on 4 real scRNA-seq datasets and 2 benchmark datasets. In these experiments, CoMem-DIPHW consistently demonstrated superior performance in differentiating and classifying distinct cell types compared to the baseline methods. Additionally, we designed a versatile simulation algorithm for scRNA-seq data that is highly customizable through its many user-specifiable parameters. This algorithm is designed to model a wide range of expression scenarios and modular structures. Notably, it incorporates intermodular co-expressions, which mimic cell-type cross-talk observed in real biological systems. For future work, we plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enhancing its scalability for larger datasets. We are also interested in exploring alternative ways to incorporate the memory mechanism into the hypergraph random walk. Currently, CoMem employs a simple product approach, moving beyond the simple product used in CoMem could potentially uncover even more effective clustering strategies.
@@
Figure/Table Description Value
Clustering results on simulated scRNA-seq data using embedding-based and community detection Methods (Figure S1) Number of cells 3000
Number of genes 1000
Number of embedded modules 30
Average number of genes per module 30
Average number of cells per module 38
Target density of the gene expression matrix 0.03
Density within modules 0.3
Density between modules 0.1
Probability of inter-module connections 0.6
Average background expression 10
Average within-module expression 20
Average inter-module expression 10
Workers 30
Walk length 30
Embedding dimensions 30
Preference exponent 50
Performance comparison of clustering methods on simulated scRNA-seq data modularity (Figure 4 for average genes per cell type plots) Average number of genes per module {10, 20, 30, 40, 70, 80, 90, 100}
Other parameters same as Figure S1
Performance comparison of clustering methods on simulated scRNA-seq data modularity (Figure 4 for number of modules plots) Number of embedded modules {10, 20, 30, 40, 70, 80, 90, 100}
Average number of genes per module 50
Average number of cells per module ( # of genes / # of modules) +3
Other parameters same as Figure S1
Clustering performance comparison by ARI across varying module sizes (Figure 5) (Same as in Figure 4 for average genes per cell type plots)
Clustering performance comparison by ARI across varying module counts (Figure 6) (Same as in Figure 4 for number of modules)
Comparison of ARI and NMI across clustering methods for the ScMixology benchmark datasets (Table 3) Number of highly expressed genes used 500
Preference exponent 10
Workers 50
Walk length 10
Embedding dimensions 10
Human pancreas (Section III.6) Number of highly expressed genes used 1000
Number of clusters 9
Preference exponent 10
Workers 100
Walk length 50
Embedding dimensions 10
Mouse pancreas (Section Additional Results for Mouse Pancreas and Mouse Brain Datasets) Number of clusters 13
Other parameters same as for human pancreas analysis (Section III.6)
Human brain (Section Additional Results for Human Brain Dataset) Number of clusters 7
Other parameters same as for human pancreas analysis (Section III.6)
Mouse brain (Section Additional Results for Mouse Pancreas and Mouse Brain Datasets) Number of clusters 9
Other parameters same as for human pancreas analysis (Section III.6)
Table S1: Hyperparameters Used in Experiments.
.1 Complexity Analysis
DIPHW
The computation cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed. Constructing the entire $|V|\times|V|$ node-to-node transition probability matrix thus involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in vectorized form is given by:
$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
While the theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the vectorized approach, it utilizes parallel computation and is significantly faster and hence more suitable for large datasets. The memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
CoMem-DIPHW
The time complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$, as for each node pair, the memory-incorporated transition traverses all possible edges that could have arrived at the first node and all possible edges that could then connect the fisrt node with the second node. The memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory incorporated node-to-edge and edge-to-node transition probability matrices.
.2 Bipartite representation
We chose the hypergraph conceptualization over bipartite representation for consistency since the DIPHW algorithm we introduced is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [41]. DIPHW extended EDVW hypergraph random walk [41] with a vertex-dependent edge selection probability, as well as a preference exponent for clustering acceleration.
A bipartite conceptualization could be adopted for our proposed methods DIPHW, CoMem, and CoMem-DIPHW by considering the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph. The two sets of nodes in the bipartite representation are the set of cells V and the set of genes E. Similarly, each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level. This alternative representation allows for the use of bipartite graph algorithms and conceptualization for analyses. For example, we used Barber’s bipartite modularity  [42] to assess the how the clustering performance of different algorithms vary with modularity of the underlying scRNA-seq dataset.
@@
.3 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
@@
There are two primary strategies for handling shared zeros in cell expression profiles of scRNA-seq data during correlation computation. The first strategy, exemplified by cosine similarity, is to ignore the shared zeros. This means that inactive expression (zeros) does not contribute to the similarity between cells. However, our experiments indicate that this strategy results in significantly worse clustering performance. Understandably, the absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded entirely.
@@
In Table S2, we evaluated how co-expression networks generated by different correlation computation methods lead to varying performance across various network-based clustering methods. The correlation methods considered were Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity. Our results show that cosine similarity often led to the worst clustering results among the correlation computation methods tested.
@@
Table S2: Impact of Correlation Methods on Network-Based Clustering Performance. The table evaluates the impact of correlation computation methods—Pearson’s, Spearman’s, Kendall’s Tau, and cosine similarity—on the performance of various network-based clustering methods. The correlation computation method resulting in the worst clustering performance is marked in bold. Cosine similarity, which ignores shared zeros in cell expression profiles, often resulted in the worst clustering performance. These results show that, although using correlation computation methods that treat zeros the same as non-zero signals could inflate correlations, ignoring these zeros, as in cosine similarity, can lead to worse performance.
@@
The primary implication of this experiment is that both strategies have inherent drawbacks when dealing with the sparsity of scRNA-seq data. Ignoring zeros misses out on significant information, while treating them as homogeneous signals inflates correlations and introduces noise. Ignoring the shared zeros in cell expression profiles, as cosine similarity does, is not an effective strategy and results in the worst performance. Conversely, methods that consider all the shared zeros in the expression profiles as an indication of cell homogeneity —such as Pearson’s, Spearman’s, and Kendall’s Tau—neglect the fact that many shared zeros are merely a result of the high dimensionality of scRNA-seq data. When sequencing results from multiple cells are merged, many zero entries are introduced since different cells often express different sets of genes. Consequently, cell correlations are inflated by these methods when data is sparse. The clustering outcomes from these methods are suboptimal when compared with approaches that do not require correlation computation, such as CoMem, DIPHW, and PCA, potentially due to misleading signals in the inflated correlation networks S1 4 5 6. Our results suggest that alternative approaches, such as hypergraph-based methods, provide a better solution by circumventing the need for unipartite network projections and avoiding the intricacy associated with handling zeros in correlation computations.
.4 Embedding Visualizations
We present the visualizations and clustering performance of embeddings generated by DIPHW, CoMem, and CoMem-DIPHW in Fig. S1. We compare these against five other embedding-based methods (PCA, Node2Vec, EDVW [41], UMAP, and t-SNE) and four community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden) applied to simulated single-cell RNA sequencing (scRNA-seq) data. Embedding-based methods compute representations of the cells in a lower-dimensional space while preserving their global geometry [77, 78, 79]. For these methods, we used UMAP for embedding visualization and K-means for clustering. In contrast, community detection methods identify clusters (communities) based on the structures of the unipartite networks projected from the scRNA-seq data. Although t-SNE and UMAP are typically not used as a standalone embedding algorithms, we have included them in the comparison for reference, as they are adopted by popular packages in single-cell analysis, such as Seurat [15, 16], Scanpy [17], SC3 [50], and Cell Ranger [26]. There has been ongoing discussions about using them as the main clustering methods in scRNA-seq analysis [80, 81]. We note that the strength of t-SNE and UMAP lies in preserving local structures, such as the nearest neighbors [80, 54], while UMAP can also preserve global distance [54]. However, they are not designed to group data into clusters with high internal cohesion. Therefore, they are generally used in conjunction with other clustering methods as visualizing tools that map computed embeddings to a 2D space for evaluation.
The results are compared with the ground-truth cell-type membership in the embedding visualization and in the NMI and ARI bar plots. For visualization, UMAP was applied to the embeddings produced by the dimensionality reduction algorithms. Each dot represents a cell, with its color indicating the ground-truth cell cluster membership and its position representing the embedding. As illustrated, DIPHW and CoMem-DIPHW demonstrated the highest degree of internal cohesion within clusters and the best cluster separation, as evidenced by both the embedding visualizations and the ARI and NMI scores. We also evaluate CoMem as a baseline method, which shares the same memory mechanism as CoMem-DIPHW but without the dual importance (edge-dependent vertex selection probability and vertex-dependent edge selection probability) or preference exponent. This comparison shows the improvements brought by integrating the hypergraph-based random walk DIPHW with the memory mechanism in CoMem.
The parameters used to simulate the scRNA-seq data in this experiment were chosen to capture the main characteristics and underlying assumptions of scRNA-seq analysis. Although these parameters may not perfectly match the values observed in real scRNA-seq data, they provide a challenging enough clustering task to differentiate the performance of the evaluated clustering algorithms. The inter-module density is set at 0.1, and inter-module connection probability at 0.6 to represent cell-type cross-talk [63, 66, 65, 64, 82], resulting in a weaker expression signal density than within cell-type expression. To distinguish cell-type-specific signals from background expression, we set the baseline gene expression level at 10 and the within-module average expression at 20.
Figure S1: Clustering Results on Simulated ScRNA-seq Data Using Embedding-based and Community Detection Methods. (a) UMAP visualization of the embedding-based clustering results, with colors indicating ground-truth cell cluster membership. (b) NMI and ARI bar plots comparing the performance of the embedding and community detection methods. Higher NMI and ARI values suggest greater consistency with ground-truth clusters, indicating better clustering performance. DIPHW and CoMem-DIPHW demonstrated the highest degree of internal cohesion within clusters and the best cluster separation, as evidenced by both the embedding visualizations and the ARI and NMI scores. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
Additional Results for Clustering Performance Across Modularity
@@
Figure S2: Clustering Performance Comparison by ARI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. Here, we present the complementary results across intermediate parameter settings, which show consistent patterns: our proposed methods (highlighted in red and yellow) consistently ranked first by ARI in all scenarios, with a stronger advantage when data modularity is weak, i.e., when the average number of co-expressed genes per module is small. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
@@
Figure S3: Clustering Performance Comparison by ARI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. Here, we present the complementary results across intermediate parameter settings, which show consistent patterns: our proposed methods (highlighted in red and yellow) consistently ranked first by ARI in all scenarios, with a stronger advantage when data modularity is weak, i.e., when the number of embedded modules in the simulated scRNA-seq data is greater. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
@@
Figure S4: Clustering Performance Comparison by NMI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. The NMI results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest NMI values or best NMI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
@@
Figure S5: Clustering Performance Comparison by NMI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. The results by NMI support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest NMI values or best NMI rankings. For all embedding-based methods, the resulting embeddings are coupled with K-means for clustering purposes.
@@
Additional Results for Human Brain Dataset

### v4

@@ IV Conclusion
In this study, we develop and analyze two hypergraph-based clustering algorithms, DIPHW and CoMem-DIPHW, to improve cell clustering performance on scRNA-seq data. First, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW) extends the edge-dependent vertex weight (EDVW) hypergraph random walk [Chitra2019] and consistently outperforms it in our experiments. The key idea of DIPHW is its symmetrized treatment of importance, considering both the relative importance of edges to nodes and nodes to edges, or the importance of genes relative to cells and cells to genes, along with a preference exponent for faster clustering. Second, the CoMem and CoMem-DIPHW clustering methods incorporate a memory mechanism that considers both the abundance counts data (conceptualized as hypergraphs) and the two co-expression networks to integrate local and global information when computing embeddings for clustering. Third, we identify and demonstrate through experiments (see Fig. 3) the issue of inflated signals in co-expression networks caused by the sparsity of scRNA-seq data, which further motivates the use of hypergraph representations. Fourth, we compare our proposed methods with 13 cell-clustering methods from various categories, including community detection, embedding-based methods, and recent deep learning methods used in clustering scRNA-seq data. Our evaluation covers a broad range of modularity conditions in the data, where CoMem-DIPHW and DIPHW consistently show superior performance, particularly in weak modularity regimes. When clusters are harder to detect, other methods experience a much more pronounced decline in performance compared to CoMem-DIPHW. Fifth, we test our proposed methods on four real scRNA-seq datasets and two benchmark datasets. In these experiments, CoMem-DIPHW consistently demonstrates superior performance in differentiating and classifying distinct cell types compared to the other methods. Sixth, we design a versatile simulation algorithm for scRNA-seq data that is highly customizable through many user-specified parameters. This algorithm models a wide range of expression scenarios and modular structures. Notably, it incorporates intermodular co-expressions, which mimic cell-type cross-talk observed in real biological systems.
IV.1 Future Work
We plan to optimize the implementation of CoMem-DIPHW to reduce its memory complexity, enabling it to scale to larger datasets. We will also explore alternative methods for incorporating the memory mechanism into the hypergraph random walk. Currently, CoMem uses a simple product approach; moving beyond this could reveal more effective clustering strategies. Although our current evaluation focuses on the biological relevance of clusters through cell-type annotation using PanglaoDB markers, future application-focused studies could include Gene Ontology enrichment analysis to characterize the functions of identified clusters.
With the increasing availability of data and studies in spatial transcriptomics [li2024spadit, rao2021exploring, longo2021integrating, kleshchevnikov2022cell2location], extending the CoMem-DIPHW formulation for spatial transcriptomics analysis is a promising direction. While the current CoMem-DIPHW formulation focuses on cell clustering based on expression profile similarity, it could be extended to incorporate spatial proximity by replacing the cell expression similarity matrix $G_{V}$ with functions that characterize joint cell similarity based on both expression profiles and spatial information.
@@ IV.1 Future Work
S2 Evaluating the Impact of Correlation Methods on Network-Based Clustering Performance
@@
There are two primary strategies for handling shared zeros in cell expression profiles of scRNA-seq data during correlation computation. The first strategy, exemplified by cosine similarity, is to ignore the shared zeros. In this approach, inactive expression (zeros) does not contribute to the similarity between cells. However, our experiments indicate that this strategy leads to significantly worse clustering performance. The absence of expression in the same genes between cell pairs provides information about cellular similarity and should not be disregarded.
@@
We evaluate how co-expression networks generated by different correlation computation methods affect the performance of various network-based clustering methods. See Table S1. Our results show that cosine similarity often produces the worst clustering results among Pearson, Spearman, and Kendall Tau.
@@
Table S1: Impact of Correlation Methods on Network-Based Clustering Performance. The table shows how different correlation measures (specifically, Pearson, Spearman, Kendall Tau, and cosine similarity) affect the performance of various network-based clustering methods. The correlation computation method that results in the worst clustering performance is marked in bold. Cosine similarity, which ignores shared zeros in cell expression profiles, often produces the worst clustering performance. These results indicate that while correlation computation methods that treat zeros the same as non-zero signals may inflate correlations, ignoring these zeros, as in cosine similarity, can lead to poorer performance.
@@
The main implication of this experiment is that both strategies have limitations when handling sparse scRNA-seq data. Cosine similarity’s omission of shared zeros results in the poorest performance. In contrast, methods that treat all shared zeros in expression profiles as indicators of cell homogeneity (such as Pearson, Spearman, and Kendall Tau) overlook the fact that many shared zeros arise simply from the high dimensionality of scRNA-seq data. When sequencing results from multiple cells are combined, many zero entries appear because different cells often express different sets of genes. Our results indicate that alternative approaches, such as hypergraph-based methods, offer a better solution by avoiding the need for unipartite network projections and the complexity of handling zeros in correlation calculations.
S3 Computational and Memory Complexities of DIPHW and CoMem-DIPHW
DIPHW.
The computational cost of each element $P(u\rightarrow v)$ in the $|V|\times|V|$ node-to-node transition probability matrix is $O(|E|)$ when the intermediate matrices for the node-to-edge and edge-to-node transition probabilities $P_{E|V}(e|u)$ and $P_{V|E}(v|e)$ are precomputed. The construction of the entire $|V|\times|V|$ node-to-node transition probability matrix involves computing these probabilities for all node pairs, resulting in an overall time complexity of $O(|V|^{2}\cdot|E|)$.
For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ in martix form is given by:
$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
The theoretical complexity remains $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$. However, the matrix representation utilizes parallel computation and is significantly faster, making it more suitable for large datasets.
DIPHW’s memory complexity is $O(|E|\cdot|V|)$ for the node-to-edge and edge-to-node transition probability matrices.
CoMem-DIPHW.
The computational complexity of CoMem-DIPHW is $O(|E|^{2}\cdot|V|^{2})$. For each pair of nodes, the memory-incorporated transition traverses all possible edges that could have reached the first node and all possible edges that could then connect the first node with the second node.
CoMem-DIPHW’s memory complexity is $O(|E|^{2}\cdot|V|+|E|\cdot|V|^{2})$ for the memory-incorporated node-to-edge and edge-to-node transition probability matrices.
S4 Bipartite representation
We choose the hypergraph conceptualization over the bipartite representation for consistency, as the DIPHW algorithm we introduce is an extension of the edge-dependent vertex weight hypergraph random walk (EDVW) [Chitra2019]. DIPHW extends the EDVW hypergraph random walk by incorporating a vertex-dependent edge selection probability and a preference exponent to accelerate clustering.
A bipartite conceptualization could also be applied to our proposed methods by interpreting the incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ as a bipartite graph. In this representation, the two sets of nodes are the set of cells $V$ and the set of genes $E$. Each gene is connected to all cells in which it is actively expressed, with weights corresponding to the expression level. This alternative representation enables the use of bipartite graph algorithms. For example, we use Barber’s bipartite modularity [barber2007modularity] to assess how the clustering performance of different algorithms varies with the modularity of the underlying scRNA-seq dataset.
S5 Quantitative Evaluation of Clustering Quality on Tissue Datasets
The quantitative evaluations in Tables S2 and S3 offer complementary perspectives on clustering performance for tissue scRNA-seq datasets. The evaluation in the learned embedding space (Table S2) reflects the geometric separability and compactness of the clusters produced by each method, while the evaluation of the preprocessed expression data (Table S3) assesses whether the same cluster assignments capture the structure of the preprocessed gene expression data used as input to the embedding methods. In the embedding space, graph-sc and scASDC achieve strong geometric separation and cohesion across datasets. However, they show reduced clustering performance when evaluated on the preprocessed data. This aligns with the results in Fig. 12, where scASDC produces embeddings with strong cluster separation and cohesion, but some identified DEGs correspond to biologically implausible cell types. In particular, PCA performs best in the preprocessed data evaluation. This is consistent with the findings of Ciortan et al. [ciortan2022gnn], where PCA outperformed many state-of-the-art methods, including graph-sc, when evaluated by the Silhouette score and the Calinski-Harabasz index on real and simulated datasets.
Human Brain Human Pancreas Mouse Brain Mouse Pancreas
Method SC CH DB CV SC CH DB CV SC CH DB CV SC CH DB CV
scASDC 0.421 244.710 0.975 2.373 0.763 7005.583 0.549 0.821 0.816 5230.004 0.640 1.845 0.874 2843.512 0.462 0.563
graph-sc 0.352 705.099 0.844 0.560 0.529 4164.981 0.584 0.816 0.418 2431.617 0.982 0.905 0.394 976.784 0.768 0.495
CoMem_DIPHW 0.285 224.443 1.207 1.287 0.384 1048.921 1.119 1.058 0.396 1222.851 0.936 1.403 0.274 313.957 1.183 1.677
CoMem 0.265 199.820 1.285 1.922 0.474 1447.208 0.852 3.575 0.455 1432.727 0.817 1.086 0.291 287.810 1.179 4.768
tsImpute 0.263 703.410 1.163 8.200 0.457 715.814 0.789 5.051 0.409 867.564 1.095 8.499 0.424 666.677 0.902 10.673
PCA 0.258 236.768 1.412 6.802 0.294 542.832 1.270 89.515 0.229 512.127 1.652 83.396 0.107 120.386 1.804 125.705
CAKE 0.156 164.736 1.679 2.849 0.417 3999.076 0.794 1.944 0.461 12194.111 0.800 2.684 0.237 373.750 1.391 1.825
node2vec 0.133 172.629 2.118 2.307 0.409 1405.125 1.034 1.359 0.332 1108.268 1.188 27.464 0.198 265.955 1.467 1.014
DIPHW 0.122 110.724 2.138 0.903 0.204 480.360 1.328 1.060 0.127 264.797 2.444 5.145 0.055 35.369 3.380 3.147
EDVW 0.109 97.634 2.315 1.720 0.152 402.944 2.039 6.156 0.126 246.265 2.616 5.840 0.060 40.552 3.482 5.804
Table S2: Clustering Quality Evaluated in the Learned Embedding Space on Tissue scRNA-seq Datasets. Clustering performance is assessed using the Silhouette score (SC), Calinski–Harabasz index (CH), Davies–Bouldin index (DB), and Coefficient of Variation (CV) on the learned embeddings produced by each method. Higher values indicate better performance for the Silhouette score and Calinski–Harabasz index, while lower values indicate better performance for the Davies–Bouldin index and Coefficient of Variation. This evaluation reflects the geometric separability and cohesion of clusters in the embedding space. The best values within each dataset/evaluation measure column are highlighted in bold, and the second-best values are underlined.
Human Brain Human Pancreas Mouse Brain Mouse Pancreas
Method SC CH DB CV SC CH DB CV SC CH DB CV SC CH DB CV
CoMem 0.027 17.529 4.654 3.951 0.071 230.969 2.048 4.559 0.041 173.331 2.838 7.603 -0.002 37.954 4.182 5.064
CoMem_DIPHW 0.027 17.529 4.575 3.951 0.021 235.958 2.364 4.834 0.018 169.640 3.363 7.795 0.002 40.049 3.891 5.008
PCA 0.025 19.080 4.991 4.012 0.276 395.673 1.461 4.012 0.087 183.428 2.829 6.995 0.016 46.001 3.051 4.699
tsImpute 0.009 12.901 7.122 3.958 0.114 171.384 2.715 3.734 -0.005 67.857 6.082 6.721 -0.025 22.239 6.907 5.127
graph-sc 0.008 10.692 6.286 4.043 0.042 201.025 2.218 4.137 0.047 127.852 5.446 6.905 0.052 28.514 3.696 4.749
node2vec 0.007 15.731 8.386 4.010 0.060 259.873 2.019 4.609 0.040 174.822 3.240 7.726 -0.001 37.939 4.459 5.027
DIPHW 0.006 14.937 7.794 4.018 0.015 195.839 3.152 4.328 0.012 153.568 4.081 7.782 -0.006 36.516 5.353 5.115
EDVW 0.002 13.243 8.843 4.042 -0.075 177.732 7.734 4.956 0.034 172.331 3.403 7.805 -0.001 39.774 4.790 4.991
scASDC -0.009 7.670 7.899 3.947 -0.128 139.358 3.605 4.655 0.070 166.681 3.032 7.177 -0.027 26.217 6.188 5.148
CAKE -0.012 7.713 8.723 4.002 -0.025 184.071 3.063 4.444 0.068 75.122 6.077 5.624 0.006 33.657 6.005 5.042
Table S3: Clustering Quality Evaluated on Preprocessed Expression Data for Tissue scRNA-seq Datasets. Using the same cluster assignments as in Supplementary Table S2, we evaluate cluster quality on the preprocessed gene expression data used as input to each embedding method. We measure cluster quality using the Silhouette score (SC), Calinski–Harabasz index (CH), Davies–Bouldin index (DB), and Coefficient of Variation (CV). Higher values indicate better performance for SC and CH, while lower values indicate better performance for DB and CV. This evaluation assesses whether the cluster assignments by each method capture meaningful structure in the input gene expression data. The best values within each dataset/evaluation measure column are highlighted in bold, and the second-best values are underlined.
S6 Additional Results for Clustering Performance Across Modularity Levels
Figures S1 and S2 show additional results on the impact of module size and module count on ARI clustering performance. Figures S3 and S4 show the impact of module size and module count on NMI clustering performance. Figures S5 and S6 show the impact of module size and module count on ACC clustering performance. Figures S7 and S8 show the impact of module size and module count on AMI clustering performance. Figures S9 and S10 show the impact of module size and module count on F1 clustering performance. Finally, Fig. S11 shows a visualization of modularity in our imulated scRNA-seq Data.
@@
Figure S1: Clustering Performance Comparison by ARI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. Here, we present the complementary results across intermediate parameter settings, which show consistent patterns: our proposed methods (highlighted in red and yellow) consistently ranked first by ARI in all scenarios, with a stronger advantage when data modularity is weak, i.e., when the average number of co-expressed genes per module is small. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
@@
Figure S2: Clustering Performance Comparison by ARI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. Here, we present the complementary results across intermediate parameter settings, which show consistent patterns: our proposed methods (highlighted in red and yellow) consistently ranked first by ARI in all scenarios, with a stronger advantage when data modularity is weak, i.e., when the number of embedded modules in the simulated scRNA-seq data is greater. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ARI values or best ARI rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
@@
Figure S3: Clustering Performance Comparison by NMI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. The NMI results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest NMI values or best NMI rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
@@
Figure S4: Clustering Performance Comparison by NMI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. The results by NMI support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest NMI values or best NMI rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure S5: Clustering Performance Comparison by ACC Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. The ACC results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ACC values or best ACC rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure S6: Clustering Performance Comparison by ACC Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. The results by ACC support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest ACC values or best ACC rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure S7: Clustering Performance Comparison by AMI Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. The AMI results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest AMI values or best AMI rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure S8: Clustering Performance Comparison by AMI Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. The results by AMI support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest AMI values or best AMI rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure S9: Clustering Performance Comparison by F1 Across Varying Module Sizes. Simulated scRNA-seq data were used for this evaluation. The F1 results support the same conclusion: when modularity is weak (i.e., when the average number of co-expressed genes per module is small), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest F1 values or best F1 rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
Figure S10: Clustering Performance Comparison by F1 Across Varying Module Counts. Simulated scRNA-seq data were used for this evaluation. The results by F1 support the same conclusion: when modularity is weak (i.e., when the number of modules is greater), the advantage of our proposed methods (highlighted in red and yellow) is more pronounced. Each experiment was repeated 10 times per parameter setting, with error bars representing the 95% confidence interval. Red dashed lines indicate the highest F1 values or best F1 rankings. K-means was used to cluster the output of all embedding-based methods that do not directly assign cluster membership.
@@
S7 Results for the Human Brain Dataset
Figures S12S13S14S15S16, and S17 show the results on the human brain dataset for CoMem-DIPHW, PCA, graph-sc, tsImpute, CAKE, and scASDC, respectively.